# Do quantum linear solvers offer advantage for networks-based system of linear equations?

### v3

@@ VII.0.2 Computations on large $A$ matrices: all-qubit fixing in special graphs
We consider two instances of the complete graph family to compute effective resistance on quantum hardware: all-to-all connected hexagon and all-to-all connected octagon. For complete graphs, $\vec{I}=\delta_{i}-\delta_{j}$ (refer section A.3) happens to be one of the eigenvectors of the Laplacian matrix. After QPE measurement, one would obtain a single peak in the histogram corresponding to the eigenvector. This, effectively, would reduce the circuit to one-qubit circuit consisting of HHL ancillary qubit. For all-to-all connected hexagon graph, we execute a one-qubit circuit containing $RY(0.3349)$ on IonQ Forte-1. The average PFD over 5 repeats with $1000$ shots was $2.72$ percent. For all-to-all connected octagon, we execute a one- qubit circuit with gate $RY(\pi)$ on IonQ Forte-1 quantum device, to obtain an average PFD of $0.200$ with $1000$ shots and $5$ repeats.
@@ VIII Conclusions
In summary, we investigate the prospects of quantum advantage using various quantum linear solvers (QLSs; see Section II) for the network based linear systems problem (NLSP; introduced in Section III) by surveying 50 graph families (random, non-random, trees, as well as those with/without sources and sinks), 30 based on Laplacian and 20 on incidence matrices. Our choice of picking NLSPs is motivated by both the flexibility that they offer in runtime complexity as well as their relevance to real-world problems. As Section IV describes, we find the scaling of condition number ($\kappa$), sparsity ($s$), and system size for each graph family, and within the scope of our survey (detailed in Section A.4), we classify each family as best, better, good, or bad, corresponding to exponential, polynomial, sub-linear, or no advantage respectively with HHL. The rationale behind picking HHL is that other QLSs from our list provided in Table 1 build on HHL. For example, a graph family which offers polynomial advantage with HHL is guaranteed to do at least as well with other considered QLSs.
Our findings on the categories of graph families are summarized below:
  1. 1.
    Only 2 of the 50 graph families considered (4 percent), both hypercube graphs, offer potential for exponential advantage. Realizing such an advantage requires that $\kappa$ and $s$ scale at most polylogarithmically in system size, and the system size itself needs to grow exponentially.
  2. 2.
    Realizing a polynomial advantage with HHL is much more common than exponential advantage; about 20 percent of the graph families belong to the better graph family category. While $\kappa$ and $s$ both grow at most polylogarithmically in all of these candidate graph families, the system size grows polynomially.
  3. 3.
    About another 18 percent graph families offer sub-linear advantage. These may not be practically useful, given other expected overheads from a fault-tolerant implementation, as emphasized in Ref. [prx2021]. In spite of $\kappa$ and $s$ both growing at most polylogarithmically in all of these graph families, the system size grows only linearly, thus limiting advantage.
  4. 4.
    We also identify graph families, trees in particular, that exhibit futile exponential advantage. where the complexity ratio indicates an exponential advantage but the QLS runtime itself is already exponential, rendering the advantage impractical. Thus, rather than rely solely on runtime ratios, one could, for a meaningful assessment of advantage, also look additionally into the absolute runtimes of QLSs.
@@
    Furthermore, the CKS(1) and the AQC(1) algorithms elevate certain graph families to classes with more advantage. In fact, 7 percent of the families classified as bad, that is, offering no advantage, under HHL show polynomial advantage with CKS(1)/AQC(1). This fraction substantially improves to about 45 percent with the fictitious dream linear solver, which in turn indicates the need to improve QLSs further to reap the benefits of better advantage prospects. The flip side is that even the dream solver leaves about 55 percent of the bad graph families either in the bad or good (graduates them to sub-linear advantage) categories, thus also indicating that there is only so much that even a dream solver can do.
Section V discusses a generalization of the hypercube and complete graph families to the generalized hypercube superfamily, which we find accommodates an infinity of best and better graph families in its tableau (see Figure 6). We anticipate that further work in the direction of finding such generalized superfamilies can aid in finding many such best and better graphs.
Lastly, we briefly recap our findings from Section VI, where we ask if one can directly look at a graph family properties and assess prospects of advantage for Laplacian matrix- based NLSPs. We examine the $A$ matrices of the 30 graph families considered in our survey. For all of them, we find the matrix elements occurring as two distinct types of patterns for two types of $\kappa$ behaviour: ‘diffuse’ and ‘sharp’ corresponding to slow (at most polylogarithmic $\kappa$ scaling) and fast (polynomial/exponential). We thus conjecture that if one finds new edges spawning from existing vertices as system size increases, one could expect slow $\kappa$ growth, and hence better prospects for advantage. One could additionally look for a combination of the graph family construction for fast system size growth and slow growing $d_{max}$ for favourable $s$ growth to comment on possible advantage expected from that graph family.
Even if a graph family were to theoretically offer exponential advantage, there are other significant practical challenges that lie on the way to truly realizing such a speed-up. We list some of the fine print in this regard (see Appendix A.6), and as proof-of-concept, perform effective resistance computations involving $(4\times 4)$ matrices on the newer generations of the commercially available IonQ quantum computers in Section VII.0.1. We find that even executing problems of this size is challenging, thus shedding light on the massive gap between dreams of achieving advantage in such classes of problems and current-day ground reality capabilities of quantum computers.
@@ Acknowledgements.
APPENDIX
Appendix A.1 Steps involved in the HHL algorithm (including feature extraction)
The HHL algorithm broadly encompasses the following steps. Detailed reviews describing the HHL algorithm in general can be found in Refs. [zaman2023step, lin2022lecture, morales]:
@@ Appendix A.1 Steps involved in the HHL algorithm (including feature extraction)
Figure A.1: Schematic of the HHL algorithm containing the HOM module to extract a feature from the solution vector.
A circuit that carries out the steps mentioned above (and for the specific example of extracting overlap between the input and output vectors) is illustrated in Fig. A.1.
Appendix A.2 Some relevant graph theory definitions
Definition A.2. 1.
@@ Appendix A.2 Some relevant graph theory definitions
$d_{ij}=\begin{cases}d(i)&\text{if}\ i=j;\\ 0&\text{otherwise}.\end{cases},$ (A.1)
@@
$q_{ij}=\begin{cases}1&\text{if}\penalty 10000\ (v_{i},v_{j})\in E;\\ 0&\text{otherwise}.\end{cases}.$ (A.2)
@@
$q_{ij}=\begin{cases}w_{ij}&\text{if}\penalty 10000\ (v_{i},v_{j})\in E;\\ 0&\text{otherwise}.\end{cases}$ (A.3)
Appendix A.3 Examples of NLSP
A.3.1  Determining effective resistance in complex networks
@@
Electrical networks can be modeled as graphs, where each resistor is replaced by an edge, with the weight of the edge $(v_{i},v_{j})\in E$ being $w_{ij}=1/r_{ij}$, where $r_{ij}$ is the resistance of the edge. Consider the graph, $G$, given in the top right panel of Fig. 1, which corresponds to an electrical network. Let $u_{1},\ u_{2},\ u_{3},\ \text{and}\ u_{4}$ represent the voltages at vertices $v_{1},\ v_{2},\ v_{3}$, and $v_{4}$ respectively. We represent the currents entering the network at vertices $v_{1},v_{2},v_{3},\ \text{and}\ v_{4}$ as $I_{1},I_{2},I_{3},\ \text{and}\ I_{4}$ respectively. We assume there is no leakage of currents in the network. Hence, the amount of current entering the vertex is equal to the amount of current exiting the vertex. We choose the direction of flow of current as $v_{1}\to v_{2}$, $v_{2}\to v_{3}$, $v_{3}\to v_{4}$ and $v_{4}\to v_{1}$. We remark here that the choice of directions are arbitrary, and it does not affect the final equations. From the conservation of currents at every vertex of the network, we have the following set of equations:
$\text{At }v_{1}:\quad$ $(w_{12}+w_{41})u_{1}-w_{12}u_{2}-w_{41}u_{4}=I_{1}$ (A.4)
$\text{At }v_{2}:\quad$ $(w_{12}+w_{23})u_{2}-w_{23}u_{3}-w_{12}u_{1}=I_{2}$ (A.5)
$\text{At }v_{3}:\quad$ $(w_{34}+w_{23})u_{3}-w_{34}u_{4}-w_{23}u_{2}=I_{3}$ (A.6)
$\text{At }v_{4}:\quad$ $(w_{41}+w_{34})u_{4}-w_{41}u_{1}-w_{34}u_{3}=I_{4}.$ (A.7)
@@
$\underbrace{\begin{bmatrix}w_{12}+w_{41}&-w_{12}&0&-w_{41}\\ -w_{12}&w_{12}+w_{23}&-w_{23}&0\\ 0&-w_{23}&w_{23}+w_{34}&-w_{34}\\ -w_{41}&0&-w_{34}&w_{34+41}\end{bmatrix}}_{\text{L}}\underbrace{\begin{bmatrix}u_{1}\\ u_{2}\\ u_{3}\\ u_{4}\end{bmatrix}}_{\vec{u}}=\underbrace{\begin{bmatrix}I_{1}\\ I_{2}\\ I_{3}\\ I_{4}\end{bmatrix}}_{\vec{I}}.$ (A.8)
@@
$L\vec{u}=\vec{I}.$ (A.9)
Here, $\vec{u}$ is the unknown vector $\vec{x}$ and $\vec{I}$ takes the place of $\vec{b}$ in Eq. 2. This equation assists us to solve for the vector of potentials. We use $\vec{u}=L^{+}\vec{I}$ to compute the effective resistance between two vertices, say $v_{1}$ and $v_{2}$ of the graph, where $L^{+}$ is pseudo-inverse of $L$. This is obtained by allowing a unit current to flow into the network at $v_{1}$ and out of the network at $v_{2}$. We define $\vec{\delta}_{i}=(0,0,\cdots,\ 1,\cdots,0)^{T}$, where $1$ occurs only at $i^{th}$ index. To find effective resistance between $v_{1}$ and $v_{2}$, we set $I_{1}=1$ unit of current and $I_{2}=-1$ unit of current in Eq. A.8. Thus, $\vec{I}=\vec{\delta}_{1}-\vec{\delta}_{2}$. As we allow only a unit current to enter and exit from vertices $v_{1}$ and $v_{2}$ respectively, finding effective resistance is as simple as finding the potential difference between vertices $v_{1}$ and $v_{2}$. Hence, $r_{\text{eff}}^{12}=u_{1}-u_{2}=(\vec{\delta}_{1}-\vec{\delta}_{2})^{T}L^{+}(\vec{\delta}_{1}-\vec{\delta}_{2})$. $\square$
A.3.2  Determining number of vehicles on a lane: traffic flow congestion detection
@@
Consider the directed graph $\overrightarrow{G}$ depicted in the top right panel of Fig. 1 where all the vertices $v_{1},v_{2},v_{3}$, and $v_{4}$ preserve the flow conservation property. Viewed as a traffic flow problem, the edges are the lanes, while the weights on the edges denote the number of vehicles on that lane. The vehicles along a lane move in a particular direction, and is given by the direction marked on an edge. The vertices are the junctions. The vertices $v_{1}$ and $v_{4}$ take $c_{1}$ and $c_{4}$ vehicles respectively into the system of lanes. On the other hand, $c_{2}$ and $c_{3}$ are the number of vehicles that go out of the system via $v_{2}$ and $v_{3}$ respectively. The values of $c_{1},c_{2},c_{3}$, and $c_{4}$ are known. Let the flows via $\overrightarrow{(v_{1},v_{2})},\overrightarrow{(v_{2},v_{3})},\overrightarrow{(v_{3},v_{4})}$ be $\overrightarrow{(v_{4},v_{1})}$ are $y_{1},y_{2},y_{3}$, and $y_{4}$, respectively, which are unknowns. As the flow conservation property holds at all the vertices, we can construct four linear equations for the vertices, which are as follows
@@
$\underbrace{\begin{pmatrix}-1&0&0&1\\ 1&-1&0&0\\ 0&1&-1&0\\ 0&0&1&-1\end{pmatrix}}_{\text{B}}\underbrace{\begin{pmatrix}y_{1}\\ y_{2}\\ y_{3}\\ y_{4}\end{pmatrix}}_{\vec{y}}=\underbrace{\begin{pmatrix}-c_{1}\\ c_{2}\\ c_{3}\\ -c_{4}\end{pmatrix}}_{\vec{c}}.$ (A.11)
The system of linear equations can be expressed as $B\vec{y}=\vec{c}\$, where $B$ is the incidence matrix of the graph $\overrightarrow{G}$, $\vec{y}$ is the vector of flows on each edge, which is unknown in the equation and $\vec{c}$ is the vector of flows entering the network at each vertex, which is a known quantity. To solve for $\vec{y}$ using a quantum algorithm, we must transform our matrix $B$ according to Eq. (4). Our new matrix equation becomes
$\underbrace{\begin{pmatrix}0&B\\ B^{\dagger}&0\end{pmatrix}}_{\text{H}}\underbrace{{\begin{pmatrix}0\\ \vec{y}\end{pmatrix}}}_{\vec{Y}}=\underbrace{{\begin{pmatrix}\vec{c}\\ 0\end{pmatrix}}}_{\vec{C}}.$ (A.12)
@@
Appendix A.4 Scope and assumptions of our numerical analyses
@@
Figure A.2: An illustration of cut-off problem with (a) ladder graph family, and (b) with ring of cliques graph family. Sub-figure (c) provides a comparison of $\tilde{R}(n)$ versus $n$ for CLS and KMP algorithms. The y-axis is on the log scale.
A.4.1  Assumptions
@@
  • $1/\epsilon$ scaling: As noted earlier, we assume for simplicity that $\epsilon^{-1}\sim\mathrm{log}(\mathcal{N})$. This assumption is reasonable, since for example, in HHL, this translates to $n_{r}$ growing as $log($$\mathcal{N}$)), where $n_{r}$ is the number of clock register qubits in the HHL quantum circuit (see Appendix A.1). This assumption can be relaxed depending on the specific application and target precision of interest in future studies, but we note that setting $\epsilon^{-1}\sim\mathcal{N}$, for example, can impact the prospects of an advantage adversely. For example, in the hypercube graph, which as we shall see in Section IV.1 is a best graph, that is, it offers an exponential advantage with HHL, such a behaviour in $\epsilon^{-1}$ leads to a loss of advantage.
@@
  • We only consider graph Laplacians with edge weight of each edge set to $1$, for simplicity. A more general version of the analysis can be performed by relaxing this constraint for a future study, but for the purposes of this work, we perform preliminary analysis to inspect the effect of this assumption in obtaining advantage. We note that changing the edge weight of 1 for all of the edges to any other fixed real scalar for all of the edges does not affect $\kappa$ or $s$. Thus, we consider three cases of non-uniform edge weights, where they grow linearly, polynomially (second order), and logarithmically, in $y$, where an edge occurs between two vertices carrying indices $x$ and $y$. We carry out the analysis on (i) the hypercube graph, and (ii) modified Margulis-Gabber-Galil graph, which offer polynomial advantage with HHL in Section IV.1. Our results are presented in Figs. A.3 and A.4. We make the important observation that edge weight values do impact prospects of advantage; the hypercube graph retains its exponential advantage with logarithmically growing edge weights, but demotes to offering no advantage at all when the edge weights grow as a linear or a polynomial (second order) function of $y$. For the weighted modified Margulis-Gabber-Galil graph family, when the edge weight grows logarithmically as well as linearly, the graph family remains in the same category, that is, it is still a better graph. If the edge weights grow polynomially (second order), we find that the graph gives a polynomial disadvantage.
@@
  • Cut-off problem: To calculate the condition number, $\kappa$, of a matrix, one must be able to distinguish the minimum eigenvalue with zero. Since we carry out numerical analyses, one can immediately see that eigenvalues smaller than a certain value may not be capturable, thus leading to an incorrect estimate for $\kappa$. For instance, if we are not able to capture the minimum eigenvalue but instead end up capturing a larger one, we underestimate $\kappa$, and thus overestimate the degree of advantage obtained from that graph family. For all of our numerical calculations, we found that a cut-off of $10^{-6}$ for minimum eigenvalue is sufficient for the $\mathcal{N}$ values up to which we calculate $\kappa$. The pertinent data, where we compare the minimum eigenvalues at a few system sizes for two different thresholds, $10^{-6}$ and $10^{-10}$, and find them to be the same, is presented in Figs. A.5 and A.6. We note that the figures are only plotted for those graphs whose minimum eigenvalues show a downward trend with $\mathcal{N}$. Furthermore, when we move to $\mathcal{N}$ values larger than those considered for our study, we find two exceptions: the ladder graph family and the ring of cliques graph family. Figs. A.2(a) and (b) highlight this ‘cut-off’ problem for these two graph families.

### v4

@@ VII.0.2 Computations on large $A$ matrices: all-qubit fixing in special graphs
We consider two instances of the complete graph family to compute effective resistance on quantum hardware: all-to-all connected hexagon and all-to-all connected octagon. For complete graphs, $\vec{I}=\delta_{i}-\delta_{j}$ (refer section S.3 of Supplemental Material) happens to be one of the eigenvectors of the Laplacian matrix. After QPE measurement, one would obtain a single peak in the histogram corresponding to the eigenvector. This, effectively, would reduce the circuit to one-qubit circuit consisting of HHL ancillary qubit. For all-to-all connected hexagon graph, we execute a one-qubit circuit containing $RY(0.3349)$ on IonQ Forte-1. The average PFD over 5 repeats with $1000$ shots was $2.72$ percent. For an all-to-all connected octagon, we execute a one- qubit circuit with gate $RY(\pi)$ on IonQ Forte-1 quantum device, to obtain an average PFD of $0.200$ with $1000$ shots and $5$ repeats.
@@ VIII Conclusions
In summary, we investigate the prospects of quantum advantage using various quantum linear solvers (QLSs; see Section II) for the networks-based linear systems problem (NLSP; introduced in Section III) by surveying 50 graph families (random, non-random, trees, as well as those with/without sources and sinks), 30 based on Laplacian and 20 on incidence matrices. Our choice of picking NLSPs is motivated by both the flexibility that they offer in runtime complexity as well as their relevance to real-world problems. As Section IV describes, we find the scaling of condition number ($\kappa$) and sparsity ($s$) for each graph family, and within the scope of our survey (detailed in Section S.4 of Supplemental Material), we classify each family as good or bad, corresponding to an advantage or no advantage respectively with HHL. The rationale behind picking HHL is that other QLSs from our list provided in Table 1 build on HHL. For example, a graph family which offers polynomial advantage with HHL is guaranteed to do at least as well with other considered QLSs.
Our findings show that 21 of the 50 graph families considered are good, that is, only they offer potential for (exponential) advantage with HHL. Realizing such an advantage almost always requires that $\kappa$ and $s$ scale at most polylogarithmically in system size (with the only exception from our dataset being the Sudoku graph family, where $\kappa\sim\sqrt{N}$). Among the 21, 7 are from Laplacian matrix based families (we considered a total of 30 of them for our study) and 14 from incidence matrix-based families (where we considered a total of 20 graph families).
@@
    Furthermore, the CKS(1) and the AQC(1) algorithms elevate several graph families to classes with more advantage. In fact, 15 of the 23 bad graph families show exponential advantage with CKS(1)/AQC(1). We find one graph family where the dream QLS offers an advantage whereas the CKS(1) and the AQC(1) approaches do not, due to a peculiar cancellation between the growth of $\kappa$ and $s$.
Section V discusses a generalization of the hypercube and complete graph families to the generalized hypercube superfamily, which we find accommodates an infinity of good graph families in its tableau (see Fig. 5). We anticipate that further work in the direction of finding such generalized superfamilies can aid in finding many such new graphs.
Lastly, we briefly recap our findings from Section VI, where we ask if one can directly look at a graph family properties and assess prospects of advantage for Laplacian matrix- based NLSPs. We examine the $A$ matrices of the 30 graph families considered in our survey. For all of them, we find the matrix elements occurring as two distinct types of patterns for two types of $\kappa$ behaviour: ‘diffuse’ and ‘sharp’ corresponding to slow (at most polylogarithmic $\kappa$ scaling) and fast (polynomial/exponential). We thus conjecture that if one finds new edges spawning from existing vertices as system size increases, one could expect slow $\kappa$ growth, and hence better prospects for advantage. One could additionally look for a graph construction that could result in slow growing $d_{max}$ for favourable $s$ growth to comment on possible advantage expected from that graph family.
Even if a graph family were to theoretically offer exponential advantage, there are other significant practical challenges that lie on the way to truly realizing such a speed-up. We list some of the fine print in this regard (see S.6 of Supplemental Material), and as proof-of-concept, perform effective resistance computations involving $(4\times 4)$ matrices on the newer generations of the commercially available IonQ quantum computers in Section VII.0.1. We find that even executing problems of this size is challenging, thus shedding light on the massive gap between dreams of achieving advantage in such classes of problems and current-day ground reality capabilities of quantum computers.
@@ Acknowledgements.
Supplemental Material
Appendix S.1 Steps involved in the HHL algorithm (including feature extraction)
The HHL algorithm [31] broadly encompasses the following steps. Detailed reviews describing the HHL algorithm in general can be found in Refs. [69, 41, 46]:
@@
Figure S.1: Schematic of the HHL algorithm containing the HOM module to extract a feature from the solution vector.
A circuit that carries out the steps mentioned above (and for the specific example of extracting overlap between the input and output vectors) is illustrated in Fig. S.1.
Appendix S.2 Some relevant graph theory definitions
Definition S.2. 1.
@@
$d_{ij}=\begin{cases}d(i)&\text{if}\ i=j;\\ 0&\text{otherwise}.\end{cases},$ (S.1)
@@
$q_{ij}=\begin{cases}1&\text{if}\penalty 10000\ (v_{i},v_{j})\in E;\\ 0&\text{otherwise}.\end{cases}.$ (S.2)
@@
$q_{ij}=\begin{cases}w_{ij}&\text{if}\penalty 10000\ (v_{i},v_{j})\in E;\\ 0&\text{otherwise}.\end{cases}$ (S.3)
Appendix S.3 Examples of NLSP
S.3.1  Determining effective resistance in complex networks
@@
Electrical networks can be modeled as graphs, where each resistor is replaced by an edge, with the weight of the edge $(v_{i},v_{j})\in E$ being $w_{ij}=1/r_{ij}$, where $r_{ij}$ is the resistance of the edge. Consider the graph, $G$, given in the top right panel of Figure 1 of the main manuscript, which corresponds to an electrical network. Let $u_{1},\ u_{2},\ u_{3},\ \text{and}\ u_{4}$ represent the voltages at vertices $v_{1},\ v_{2},\ v_{3}$, and $v_{4}$ respectively. We represent the currents entering the network at vertices $v_{1},v_{2},v_{3},\ \text{and}\ v_{4}$ as $I_{1},I_{2},I_{3},\ \text{and}\ I_{4}$ respectively. We assume there is no leakage of currents in the network. Hence, the amount of current entering the vertex is equal to the amount of current exiting the vertex. We choose the direction of flow of current as $v_{1}\to v_{2}$, $v_{2}\to v_{3}$, $v_{3}\to v_{4}$ and $v_{4}\to v_{1}$. We remark here that the choice of directions are arbitrary, and it does not affect the final equations. From the conservation of currents at every vertex of the network, we have the following set of equations:
$\text{At }v_{1}:\quad$ $(w_{12}+w_{41})u_{1}-w_{12}u_{2}-w_{41}u_{4}=I_{1}$ (S.4)
$\text{At }v_{2}:\quad$ $(w_{12}+w_{23})u_{2}-w_{23}u_{3}-w_{12}u_{1}=I_{2}$ (S.5)
$\text{At }v_{3}:\quad$ $(w_{34}+w_{23})u_{3}-w_{34}u_{4}-w_{23}u_{2}=I_{3}$ (S.6)
$\text{At }v_{4}:\quad$ $(w_{41}+w_{34})u_{4}-w_{41}u_{1}-w_{34}u_{3}=I_{4}.$ (S.7)
@@
$\underbrace{\begin{bmatrix}w_{12}+w_{41}&-w_{12}&0&-w_{41}\\ -w_{12}&w_{12}+w_{23}&-w_{23}&0\\ 0&-w_{23}&w_{23}+w_{34}&-w_{34}\\ -w_{41}&0&-w_{34}&w_{34+41}\end{bmatrix}}_{\text{L}}\underbrace{\begin{bmatrix}u_{1}\\ u_{2}\\ u_{3}\\ u_{4}\end{bmatrix}}_{\vec{u}}=\underbrace{\begin{bmatrix}I_{1}\\ I_{2}\\ I_{3}\\ I_{4}\end{bmatrix}}_{\vec{I}}.$ (S.8)
@@
$L\vec{u}=\vec{I}.$ (S.9)
Here, $\vec{u}$ is the unknown vector $\vec{x}$ and $\vec{I}$ takes the place of $\vec{b}$ in equation $L\vec{x}=\vec{b}$. This equation assists us to solve for the vector of potentials. We use $\vec{u}=L^{+}\vec{I}$ to compute the effective resistance between two vertices, say $v_{1}$ and $v_{2}$ of the graph, where $L^{+}$ is pseudo-inverse of $L$. This is obtained by allowing a unit current to flow into the network at $v_{1}$ and out of the network at $v_{2}$. We define $\vec{\delta}_{i}=(0,0,\cdots,\ 1,\cdots,0)^{T}$, where $1$ occurs only at $i^{th}$ index. To find effective resistance between $v_{1}$ and $v_{2}$, we set $I_{1}=1$ unit of current and $I_{2}=-1$ unit of current in Eq. S.8. Thus, $\vec{I}=\vec{\delta}_{1}-\vec{\delta}_{2}$. As we allow only a unit current to enter and exit from vertices $v_{1}$ and $v_{2}$ respectively, finding effective resistance is as simple as finding the potential difference between vertices $v_{1}$ and $v_{2}$. Hence, $r_{\text{eff}}^{12}=u_{1}-u_{2}=(\vec{\delta}_{1}-\vec{\delta}_{2})^{T}L^{+}(\vec{\delta}_{1}-\vec{\delta}_{2})$. $\square$
S.3.2  Determining number of vehicles on a lane: traffic flow congestion detection
@@
Consider the directed graph $\overrightarrow{G}$ depicted in the top right panel of Figure 1 of the main manuscript where all the vertices $v_{1},v_{2},v_{3}$, and $v_{4}$ preserve the flow conservation property. Viewed as a traffic flow problem, the edges are the lanes, while the weights on the edges denote the number of vehicles on that lane. The vehicles along a lane move in a particular direction, and is given by the direction marked on an edge. The vertices are the junctions. The vertices $v_{1}$ and $v_{4}$ take $c_{1}$ and $c_{4}$ vehicles respectively into the system of lanes. On the other hand, $c_{2}$ and $c_{3}$ are the number of vehicles that go out of the system via $v_{2}$ and $v_{3}$ respectively. The values of $c_{1},c_{2},c_{3}$, and $c_{4}$ are known. Let the flows via $\overrightarrow{(v_{1},v_{2})},\overrightarrow{(v_{2},v_{3})},\overrightarrow{(v_{3},v_{4})}$ be $\overrightarrow{(v_{4},v_{1})}$ are $y_{1},y_{2},y_{3}$, and $y_{4}$, respectively, which are unknowns. As the flow conservation property holds at all the vertices, we can construct four linear equations for the vertices, which are as follows
@@
$\underbrace{\begin{pmatrix}-1&0&0&1\\ 1&-1&0&0\\ 0&1&-1&0\\ 0&0&1&-1\end{pmatrix}}_{\text{B}}\underbrace{\begin{pmatrix}y_{1}\\ y_{2}\\ y_{3}\\ y_{4}\end{pmatrix}}_{\vec{y}}=\underbrace{\begin{pmatrix}-c_{1}\\ c_{2}\\ c_{3}\\ -c_{4}\end{pmatrix}}_{\vec{c}}.$ (S.11)
The system of linear equations can be expressed as $B\vec{y}=\vec{c}\$, where $B$ is the incidence matrix of the graph $\overrightarrow{G}$, $\vec{y}$ is the vector of flows on each edge, which is unknown in the equation and $\vec{c}$ is the vector of flows entering the network at each vertex, which is a known quantity. To solve for $\vec{y}$ using a quantum algorithm, we must transform our matrix $B$ according to Equation (6) of the main manuscript. Our new matrix equation becomes
$\underbrace{\begin{pmatrix}0&B\\ B^{\dagger}&0\end{pmatrix}}_{\text{H}}\underbrace{{\begin{pmatrix}0\\ \vec{y}\end{pmatrix}}}_{\vec{Y}}=\underbrace{{\begin{pmatrix}\vec{c}\\ 0\end{pmatrix}}}_{\vec{C}}.$ (S.12)
@@
Appendix S.4 Scope and assumptions of our numerical analyses
@@
Figure S.2: An illustration of cut-off problem with (a) ladder graph family, and (b) with ring of cliques graph family. Sub-figure (c) provides a comparison of $\tilde{R}(n)$ versus $n$ for CLS and KMP algorithms. The y-axis is on the log scale.
S.4.1  Assumptions
@@
  • $1/\epsilon$ scaling: As noted earlier, we assume for simplicity that $\epsilon^{-1}\sim\mathrm{log}(\mathcal{N})$. This assumption is reasonable, since for example, in HHL, this translates to $n_{r}$ growing as $\log(\log(\mathcal{N}))$, where $n_{r}$ is the number of clock register qubits in the HHL quantum circuit Section S.1). This assumption can be relaxed depending on the specific application and target precision of interest in future studies.
@@
  • We only consider graph Laplacians with edge weight of each edge set to $1$, for simplicity. A more general version of the analysis can be performed by relaxing this constraint for a future study, but for the purposes of this work, we perform preliminary analysis to inspect the effect of this assumption in obtaining advantage. We note that changing the edge weight of 1 for all of the edges to any other fixed real scalar for all of the edges does not affect $\kappa$ or $s$. Thus, we consider three cases of non-uniform edge weights, where they grow linearly, polynomially (second order), and logarithmically, in $y$, where an edge occurs between two vertices carrying indices $x$ and $y$. We carry out the analysis on two of the good graphs (refer Table II of the main manuscript): (i) the hypercube graph, and (ii) modified Margulis-Gabber-Galil graph. Our results are presented in Figs. S.3 and S.4. We make the important observation that edge weight values do impact prospects of advantage; the hypercube graph retains its exponential advantage with HHL, CKS(1)/AQC(1) and dream QLS (refer Table I of main manuscript) for logarithmically growing edge weights, but becomes a bad graph (i.e., no advantage with HHL) when edge weights grow as a linear or a polynomial (second order) function of $y$. However, we do get exponential advantage with CKS(1)/AQC(1) for linearly or polynomially growing edge weights, while these scenarios show no advantage with dream QLS. For weighted modified Margulis-Gabber-Galil graph family, when the edge weight grows logarithmically as well as linearly, the graph family still remains in the good category and exhibits exponential advantage for HHL, CKS(1)/AQC(1) and the dream QLS. If the the edge weights grow polynomially (second order), the graph family is demoted to a bad graph family, which implies no advantage with HHL algorithm. We also find that, with polynomially growing edge weights, CKS(1)/AQC(1) and dream QLS show exponential advantage.
@@
  • Cut-off problem: To calculate the condition number, $\kappa$, of a matrix, one must be able to distinguish the minimum eigenvalue with zero. Since we carry out numerical analyses, one can immediately see that eigenvalues smaller than a certain value may not be capturable, thus leading to an incorrect estimate for $\kappa$. For instance, if we are not able to capture the minimum eigenvalue but instead end up capturing a larger one, we underestimate $\kappa$, and thus overestimate the degree of advantage obtained from that graph family. For all of our numerical calculations, we found that a cut-off of $10^{-6}$ for minimum eigenvalue is sufficient for the $\mathcal{N}$ values up to which we calculate $\kappa$. The pertinent data, where we compare the minimum eigenvalues at a few system sizes for two different thresholds, $10^{-6}$ and $10^{-10}$, and find them to be the same, is presented in Figs. S.5 and S.6. We note that the figures are only plotted for those graphs whose minimum eigenvalues show a downward trend with $\mathcal{N}$. Furthermore, when we move to $\mathcal{N}$ values larger than those considered for our study, we find two exceptions: the ladder graph family and the ring of cliques graph family. Figs. S.2(a) and (b) highlight this ‘cut-off’ problem for these two graph families.